OpenStudy (anonymous):

use the quadratic formula to solve the equation -x^2+6x-5=0 A. -5,-1 B. 1,5 C. -5,11 D. 2,10

OpenStudy (haseeb96):

a=-2, b= +6 , c=-5 by using quadratic formula \[\frac{ -b \pm \sqrt{b^{2}-4ac} }{ 2a }\] put the values \[\frac{ -6\pm \sqrt{-6^{2}-4(-2)(-5)} }{ 2(-2) }\] \[\frac{ -6\pm \sqrt{36-40} }{ -4 }\] \[\frac{ -6\pm \sqrt{-4} }{ -4 }\] \[\frac{ -6\pm 2i }{ -4 }\]

OpenStudy (haseeb96):

-6+2i\-4 OR -6-2i\-4 -2taking common from both equations \[\frac{ -2(3-i) }{ -4 } OR \frac{ -2(3+i) }{ -4 }\] \[\frac{ 3-i }{ 2 } OR \frac{ 3+i }{ 2 }\] this is the answer

OpenStudy (phi):

Haseeb96, a= -1 (not -2) so your answer is not quite right

OpenStudy (jack1):

@rockstargirl17 ... hust use the equation hasseeb posted, but ust the correct values for a b and c y = -x^2+6x-5 = 0 ax^2 + bx + c so a = -1 b = 6 c = -5

OpenStudy (jack1):

*just use ... hust isn't a word ... yet

OpenStudy (anonymous):

where do you plug it in to though im so confused

OpenStudy (jack1):

quadratic formula: 0 = -b + or - sqrt (b^2 - 4*a*c) / 2a so looks like:\[0 = \frac {-b \pm \sqrt (b^2 - 4 \times a \times c)}{ 2a}\]

OpenStudy (anonymous):

so it would be 0=-6+squ(6^2-4x-1x-5/ 2(-1) right?

OpenStudy (jack1):

yep, that's a plus or minus sign tho, so you need to do both so 0=-6+squ(6^2-4x-1x-5/ 2(-1) or 0=-6-squ(6^2-4x-1x-5/ 2(-1) and your using the x as times... right? (not as X)

OpenStudy (anonymous):

yes im using it as times

OpenStudy (jack1):

sweet, so work out both equations, and you'll get 2 different values. those values are x values for when y = 0 those are your answers

OpenStudy (anonymous):

so like do you just do the top first or the bottom and how would you get to answers

OpenStudy (austinl):

We have, \(\displaystyle -x^2+6x-5=0\Leftrightarrow-ax^2+bx-c=0\) \(a=-1\\ b=6\\ c=-5\) The quadratic formula is, \(\displaystyle x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) So therefore we have, \(\displaystyle x=\frac{-6\pm\sqrt{6^2-4(-1)(-5)}}{2(-1)}=\frac{-6\pm\sqrt{36-20}}{-2}=\frac{-6\pm\sqrt{16}}{-2}=\) \(\displaystyle x=\frac{-6\pm4}{-2} \) We should get 2 values for x.

OpenStudy (jack1):

so it's like: y = 0 @ x = -6+squ(6^2-4x-1x-5/ 2(-1) = -6 + squ(36 - 20)/ -2 = -6 + sqrt 16 / -2 = -6 + 4 / -2 = -2/-2 = 1 so y = 0 @ x = 1 that's one point (1, 0) now find the other: y = 0 @ x = -6 - squ(6^2-4x-1x-5/ 2(-1) (note the minus sign)

OpenStudy (jack1):

so can u solve the 2nd x value @rockstargirl17 ?

OpenStudy (austinl):

If my notation was confusing... \(\displaystyle x=\frac{-6\pm4}{-2}\) \(\displaystyle~~\rightarrow\frac{-6+4}{-2}=\frac{-2}{-2}=1\) \(\displaystyle~~\rightarrow\frac{-6-4}{-2}\frac{-10}{-2}=~?\)

OpenStudy (anonymous):

okay so all together i got sq(160)/-2 12.64/-2 = i dont think im doing it right

OpenStudy (jack1):

hmm... should be: y = 0 @ x = { -6 - squ(6^2 - 4x-1x-5 }/ 2(-1) = { -6 - squ(36 - 20) }/ -2 = {-6 - sqrt 16} / -2 = {-6 - 4} / -2 = -10/-2 x = 5 so 2nd point is y = 0 at x = 5 (5, 0)

OpenStudy (austinl):

@rockstargirl17 did you look at the work that I posted up above. That should tell you where you went wrong in your application of the quadratic formula. Compare your work to mine.

OpenStudy (jack1):

not sure where u were getting the 160 from, sorry @austinL 's instructions are a bit clearer if that helps?

OpenStudy (anonymous):

so its 5,1 but wouldnt it be -5,-1 ?

OpenStudy (jack1):

nope, negative divided by a negative = positive

OpenStudy (anonymous):

yeah but the answer choices say -5,-1 and i got the same thing

OpenStudy (austinl):

Not sure how they got that.... \(\displaystyle \frac{-a}{-a}=1\) \(\displaystyle \frac{-a}{-b}=\frac{a}{b}\)

OpenStudy (jack1):